OpenStudy (traylea):

Write an equation of the line that passes through (3,5) and is perpendicular to the graph of y=-3x+7. Write your final equation in slope-intercept form.

OpenStudy (mathmale):

Two lines with slopes equal to \[m _{1}=\frac{ -1 }{ m _{2} }\] respectively are perpendicular to each other, and vice versa. What is the slope of a line perpendicular to y=-3x+7?

OpenStudy (mathmale):

Let the slope of the given line, y=-3x+7, be \[m _{2}.\]

OpenStudy (mathmale):

Find the slope of a line perpendicular to y=-3x+7 using the relationship between slopes that I've given you.

OpenStudy (traylea):

m2=-2?

OpenStudy (saov):

Remember the slope of a rect, is the coeficient of x variable (independt variable)

OpenStudy (mathmale):

Let the slope of the given line, y=-3x+7, be m_2. I don't see where your "-2" came from. Try again. What is the slope of the given line, m_2?

OpenStudy (saov):

the slope of a rect, is the coeficient of x variable, but it is only true for a rect in the slope-intercept form. For example: y=4x -4 , in this case the slope is m = 4

OpenStudy (mathmale):

So, in y=-3x+7, the slope is what? m=?

OpenStudy (mathmale):

Compare the following: y=-3x+7 y=mx+b m=? b=?

OpenStudy (traylea):

sorry the slope would be 3. m=3 b=7 right? @mathmale

OpenStudy (mathmale):

Take a closer look at y=-3x+7. What's the slope? It's not 3.

OpenStudy (traylea):

7? @mathmale

OpenStudy (mathmale):

I asked you to compare: y=-3x + 7 y= mx + b y compares to y, -3 compares to m b compares to 7 What do "m" and "b" mean in this context (the equation of a straight line)?

OpenStudy (traylea):

the starting and ending coordinates?

OpenStudy (mathmale):

Sounds like you need to do some basic prep (reading). y=mx+b is the most common form of the equation of a straight line. It's a "linear function" because the graph has the form of a straight line. m represents the slope of the line and is equal to (rise) / (run) , or (change in y) / (change in x). b represents the y-intercept and is the y-coordinate of that point; the line crosses the y-axis at (0,b). Have you learned this before?

OpenStudy (traylea):

We've barely gone over it, my math teacher doesn't seem to care much on going over a topic if you don't understand it. So we touched base on it but it's to the point where I don't understand it at all and he's giving me things expecting me to know how to do it when I've said I didn't know how.

OpenStudy (traylea):

I know I'm a difficult student when it comes to things like this so it'll be perfectly fine with me if you'd like to move on, I can go read some other things about it and see if maybe that would help me.

OpenStudy (mathmale):

sorry about your situation; sounds difficult and frustrating. You are studying "functions," and "straight lines" or "linear functions" are in that general topic ("functions"). Is the word "function" familiar to you in the context of math?

OpenStudy (traylea):

Yes

OpenStudy (mathmale):

The equations of linear functions typically look like y=mx+b (which we've been discussing). Quadratic functions typically look like y=ax^2 + bx + c. Cubic functions: y=ax^3 + bx^2 + cx + d. All you need to know and understand at this point are the meanings of the words/ concepts "function" and "linear functins." Any questions so far?

OpenStudy (mathmale):

What does "y=mx+b" mean to you at this point?

OpenStudy (mathmale):

Have you an algebra textbook or other printed learning materials? Have you online learnign materials? Which approaches to learning seem to work best for you? What approaches to helping (by another person) seem to work best for you?

OpenStudy (traylea):

I'm so sorry for no responding, I got a call. I haven't learned on any materials really and working through a problem with me definitely seems to help me, like showing the answer and explaining how they got the answer by working through it and explaining it very thoroughly. @mathmale

OpenStudy (traylea):

I'm definitely in a frustrating situation when it comes to how I learn and I'm looking change schools next year so I don't have to go through this because it honestly makes me feel quite dumb.

OpenStudy (traylea):

y=1/3x+7

OpenStudy (mathmale):

Is there any chance you could find a tutor through your school? Does your school have an academic support department? Sounds as tho' you'd benefit from caring, ongoing support as you try to learn algebra.

OpenStudy (traylea):

I think there might be, I'll have to look into it.

OpenStudy (mathmale):

Time is the issue here. You'll need to identify concepts and vocab that you DO already understand and make a similar list for those you DO NOT understand. Concentrate on reviewing the former and learning the latter. Your original question asks that you write the equation of a line that passes thru a given point and is perpendicular to a given line. This assumes that you understand how the slopes of parallel and perpendicular lines are related, as well as understand that there are several forms of equations of straight lines. Math builds heavily upon topics previously learned, so somehow you'll need to look for connections between old and new material as you move forward.

OpenStudy (traylea):

Okay, thank you.